a storekeeper has cashew nuts that sell for $130 per kilogram and peanuts that sell for $60 per kilogram. He wishes to make a mixture of 50kg of the two kinds of nuts so that he can sell it for $78 per kilogram. how many kilograms of each will he have to put into the mixture?
Answer by lwsshak3(11628)

a storekeeper has cashew nuts that sell for $130 per kilogram and peanuts that sell for $60 per kilogram. He wishes to make a mixture of 50kg of the two kinds of nuts so that he can sell it for $78 per kilogram. how many kilograms of each will he have to put into the mixture?
..
let x=kg of cashew nuts
50-x= kg of peanuts
..
130x+60(50-x)=78(50)
130x+3000-60x=3900
70x=900
x=12.86
50-x=50-12.86=37.14
ans:
mixture:
12.86 kg of cashew nuts
37.14 kg of peanuts